building A is 150 feet shorter than building B. The total height of the two building is 1490 feet. Find the height of each building
Answer by best82(13)

In this case, you would write an equation. The equation must have one variable, not two.
So, the equation would be: 1490=B+(B-150).
1490= 2b-150
1490-150=2b
1340=2b
1340/2=b
670=b
Now see if building a is 670 feet.
670+(670+150)=1490?
670+820=1490
So,
Building A is 670 feet, and Building B is 820 feet.